# USTC-ICTS/PCFT-25-44

### v2

@@ 4.2 Euclidean path integral near extremality and zero modes
In addition to the graviton zero modes, we also need to consider the contribution from the photon zero modes due to the presence of the gauge field (4.2). Since we are using the Lorentz gauge, the determination of the photon zero modes follows the same procedure in subsection 3.2. For the background extremal geometry (4.18), the orthonormalized photon zero modes can be written as
@@ 4.3 Quantum corrections to the entropy
Finite quantum corrections to the partition function are obtained by adding the finite temperature corrections (4.20) and (4.26) to their extremal configurations (4.18) and (4.19) in the path integral over the zero modes (4.32) and (4.33). The corrections to the eigenvalues of the graviton and photon zero modes are given by the formulae (3.53) and (3.55), respectively. After substituting the zero modes and the near extremal configurations into (3.53) and (3.55) and retaining all the $\mathcal{O}(T)$ terms, the integrations yield
@@
Accordingly, the thermal partition function of the accelerating Kerr-Newmann black hole acquires low temperature quantum corrections of the form
@@ 5 Summary and discussion
In this paper, we systematically study the low temperature quantum corrections to the thermodynamics of the near extremal rotating and charged black holes with acceleration in four dimensions. The quantum corrections originate from the zero modes, which are the residual gauge fluctuations of the fields under gauge fixing, in the Euclidean path integral. Depending on the near horizon extremal geometry and the field contents, the zero modes have different behaviours. The metric fluctuations under large diffeomorphisms are known as the gravitational zero modes and the pure gauge fluctuations of the gauge field are specified as the photon zero modes. Each class of the zero modes is found to contribution a universal logarithmic term to the black hole thermal entropy at low temperature.
@@
We first consider the accelerating Kerr black hole. As discussed in section 2, the thermodynamic quantities defined by using semiclassical analysis exhibit problematic low temperature behaviours, since the black hole has no adequate thermal energy to radiate even a single Hawking quantum once its temperature drops to sufficiently small values. The way to resolve the inconsistency between the thermal energy and the Hawking radiation at low temperature is to incorporate quantum corrections to the black hole thermal partition function. Among all the fluctuations in the Euclidean path integral of the partition function, zero modes are found to contribute leading quantum corrections at low temperature. Invoking the harmonic gauge, the zero modes in the accelerating Kerr black hole case are those metric fluctuations on the extremal background that satisfy the transverse and traceless conditions. At low temperature, the Euclidean path integral of these zero modes produce a $(3/2)\log T$ term correction to the black hole entropy and a $(3/2)T$ term correction to the black hole mass. The critical temperature below which these quantum corrections become important is found to depend on the extremal horizon radius as well as the acceleration of the black hole. For fixed horizon radius, the critical temperature gets higher as the acceleration of the black hole increases.
@@
We next consider the accelerating Reissner-Nordström black hole case. As presented in section 3, the small temperature expansion of the black hole mass from semiclassical analysis still points to the mass gap puzzle. Compared to the neutral rotating case, the zero modes structure becomes richer in this charged static case. The near horizon extremal geometry is a direct product of AdS2 and S2 with warping factors. This structure consequently enables us to divide the large diffeomorphisms into two linearly independent classes: one consists of the diffeomorphisms on the AdS2 which generate the gravitational zero modes of tensor type, and the other includes the deformations of the S2 along the AdS2 which correspond to the gravitational zero modes of vector type. It is worth to note that in the absence of acceleration, the Reissner-Nordström black hole has three sets of vector modes since there are three independent Killing vectors on the sphere. The presence of acceleration kills two of them with only translation along the polar angle remaining. So there is only one set of the vector modes in the accelerating case. At low temperature, the path integral of the tensor modes produce a $(3/2)\log T$ term correction to the black hole entropy while the vector modes contribute a $(1/2)\log T$ term correction to the entropy. In addition, we also have the photon zero modes from the gauge field, which are the large gauge transformations under the Lorentz gauge fixing. The path integral of the photon zero modes gives another $(1/2)\log T$ term correction to the black hole entropy. The total contribution from the zero modes to the black hole entropy in the accelerating Reissner-Nordström case at low temperature is therefore $(5/2)\log T$.
@@
We finally discuss the low temperature quantum corrections to the accelerating Kerr-Newmann black hole in section 4. Its near horizon extremal geometry shares a similar structure to the accelerating Kerr case, which is a warped and twist product of AdS2 and S2. This determines that the gravitational zero modes are purely tensor type. The path integral of the gravitational tensor modes therefore contributes a $(3/2)\log T$ term correction to the black hole entropy at low temperature. Since this is a charged black hole, the gauge field contains also the photon zero modes, whose path integral leads to a $(1/2)\log T$ term correction in the low temperature expansion of the entropy. So the total contribution from the zero modes to the black hole entropy in the accelerating Kerr-Newmann case at low temperature is $2\log T$.
@@
To sum up, we worked out the universal $\log T$ corrections to the low temperature thermodynamics of accelerating black holes. The results are similar to those in the non-accelerating cases like Kerr and Reissner-Nordström black holes. The acceleration affects the near horizon extremal geometry with warping factors and therefore changes the number of independent Killing vectors on the deformed sphere. This in turn changes the number of sets of independent gravitation vector modes which are induced by the large diffeomorphisms on the deformed sphere.
Acknowledgement
We are grateful to Jianxin Lu, Jun Nian and Jingchao Zhang for helpful discussions. This work is supported by the National Natural Science Foundation of China under Grant No. 12105045 and No. 12247103.

### v3

@@ 4.2 Euclidean path integral near extremality and zero modes
In addition to the graviton zero modes, we also need to consider the contribution from the photon zero modes due to the presence of the gauge field (4.2). Since we are using the Lorentz gauge, the determination of the photon zero modes follows the same procedure in Sec. 3.2. For the background extremal geometry (4.18), the orthonormalized photon zero modes can be written as
@@ 4.3 Quantum corrections to the entropy
Finite quantum corrections to the partition function are obtained by adding the finite temperature corrections (4.20) and (4.26) to their extremal configurations (4.18) and (4.19) in the path integral over the zero modes (4.32) and (4.33). The corrections to the eigenvalues of the graviton and photon zero modes are given by the formulas (3.53) and (3.55), respectively. After substituting the zero modes and the near extremal configurations into (3.53) and (3.55) and retaining all the $\mathcal{O}(T)$ terms, the integrations yield
@@
Accordingly, the thermal partition function of the accelerating Kerr-Newmann black hole acquires low-temperature quantum corrections of the form
@@ 5 Summary and discussion
In this paper, we systematically study the low-temperature quantum corrections to the thermodynamics of the near extremal rotating and charged black holes with acceleration in four dimensions. The quantum corrections originate from the zero modes, which are the residual gauge fluctuations of the fields under gauge fixing, in the Euclidean path integral. Depending on the near horizon extremal geometry and the field contents, the zero modes have different behaviors. The metric fluctuations under large diffeomorphisms are known as the gravitational zero modes and the pure gauge fluctuations of the gauge field are specified as the photon zero modes. Each class of the zero modes is found to contribute a universal logarithmic term to the black hole thermal entropy at low temperature.
@@
We first consider the accelerating Kerr black hole. As discussed in Sec. 2, the thermodynamic quantities defined by using semiclassical analysis exhibit problematic low-temperature behaviors, since the black hole has no adequate thermal energy to radiate even a single Hawking quantum once its temperature drops to sufficiently small values. The way to resolve the inconsistency between the thermal energy and the Hawking radiation at low temperature is to incorporate quantum corrections to the black hole thermal partition function. Among all the fluctuations in the Euclidean path integral of the partition function, zero modes are found to contribute leading quantum corrections at low temperature. Invoking the harmonic gauge, the zero modes in the accelerating Kerr black hole case are those metric fluctuations on the extremal background that satisfy the transverse and traceless conditions. At low temperature, the Euclidean path integral of these zero modes produce a $(3/2)\log T$ term correction to the black hole entropy and a $(3/2)T$ term correction to the black hole mass. The critical temperature below which these quantum corrections become important is found to depend on the extremal horizon radius as well as the acceleration of the black hole. For fixed horizon radius, the critical temperature gets higher as the acceleration of the black hole increases.
@@
We next consider the accelerating Reissner-Nordström black hole case. As presented in Sec. 3, the small temperature expansion of the black hole mass from semiclassical analysis still points to the mass gap puzzle. Compared to the neutral rotating case, the zero modes structure becomes richer in this charged static case. The near horizon extremal geometry is a direct product of AdS2 and S2 with warping factors. This structure consequently enables us to divide the large diffeomorphisms into two linearly independent classes: one consists of the diffeomorphisms on the AdS2 which generate the gravitational zero modes of tensor type, and the other includes the deformations of the S2 along the AdS2 which correspond to the gravitational zero modes of vector type. It is worth noting that in the absence of acceleration, the Reissner-Nordström black hole has three sets of vector modes since there are three independent Killing vectors on the sphere. The presence of acceleration kills two of them with only translation along the polar angle remaining. So there is only one set of the vector modes in the accelerating case. At low temperature, the path integral of the tensor modes produce a $(3/2)\log T$ term correction to the black hole entropy while the vector modes contribute a $(1/2)\log T$ term correction to the entropy. In addition, we also have the photon zero modes from the gauge field, which are the large gauge transformations under the Lorentz gauge fixing. The path integral of the photon zero modes gives another $(1/2)\log T$ term correction to the black hole entropy. The total contribution from the zero modes to the black hole entropy in the accelerating Reissner-Nordström case at low temperature is therefore $(5/2)\log T$.
@@
We finally discuss the low-temperature quantum corrections to the accelerating Kerr-Newmann black hole in Sec. 4. Its near horizon extremal geometry shares a similar structure to the accelerating Kerr case, which is a warped and twisted product of AdS2 and S2. This determines that the gravitational zero modes are purely tensor type. The path integral of the gravitational tensor modes therefore contributes a $(3/2)\log T$ term correction to the black hole entropy at low temperature. Since this is a charged black hole, the gauge field contains also the photon zero modes, whose path integral leads to a $(1/2)\log T$ term correction in the low-temperature expansion of the entropy. So the total contribution from the zero modes to the black hole entropy in the accelerating Kerr-Newmann case at low temperature is $2\log T$.
@@
To sum up, we worked out the universal $\log T$ corrections to the low-temperature thermodynamics of accelerating black holes. The results are similar to those in the nonaccelerating cases like Kerr and Reissner-Nordström black holes. The acceleration affects the near horizon extremal geometry with warping factors and therefore changes the number of independent Killing vectors on the deformed sphere. This in turn changes the number of sets of independent gravitation vector modes which are induced by the large diffeomorphisms on the deformed sphere.
Acknowledgment
We are grateful to Jianxin Lu, Jun Nian and Jingchao Zhang for helpful discussions. This work is supported by the National Natural Science Foundation of China under Grants No. 12105045 and No. 12247103.